OpenStudy (anonymous):

Emily has 3 yellow, 8 green, 5 red, and 14 blue ankle bracelets in a bag. If Emily randomly selects an ankle bracelet out of the bag, what is the probability it will be green or blue?

OpenStudy (anonymous):

there are 3+8+5+14=30 bracelets in the bag. i can choose from 8 green and 14 blue = 22 choices. i think the answer is 22/30

OpenStudy (anonymous):

22/30

OpenStudy (anonymous):

or 11/15 if you simplify

OpenStudy (anonymous):

the probability of green being selected = 8/30 the probability of blue being selected = 14/30 therefore the probability of blue OR green being selected is 8/30 + 14/30 = 22/30 = 11/15